In a copper wire of diameter 2mm, 3×1023 electrons flow in one second. Calculate the current






1
Expert's answer
2022-02-28T13:52:30-0500

The electrical current

"I=\frac{q}{t}=\frac{Ne}{t}\\
 =\frac{3*10^{23}*1.6*10^{-19}}{1.0}=48000\:\rm A"